HUBBLE SPACE TELESCOPE DAILY REPORT #1770 PERIOD COVERED: 0000Z (UTC) 12/02/96 - 0000Z (UTC) 12/03/96 Daily Status Report as of 338/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6119 (Snapshot Survey For Companions of = Planetary Nebula Nuclei) The WF/PC-2 was used to make images of the planetary nebui = NGC-4361 in order to search for companions to its nuclei. The = observations were executed as scheduled, and no anomalies were noted. 1.2 Completed FOS/RD 6577 (An Expanded Survey for Study of the = Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red Side) was used to make = observations of the QSO UM485 in order to make a survey of the Lyman-Alpha= line and the 68 Mg II absorption line system. The observations were = completed as scheduled, and no problems were reported. 1.3 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.4 Completed WF/PC-2 6344 (Low-Mass Binaries in the Hyades - = Completing the Survey) The WF/PC-2 was used to make observations of the low-mass binary = star RHY406 (M-Class Dwarf Stars) which is located in the Hyades Cluster. = The observations were executed as scheduled, and no anomalies were = noted. 1.5 Completed S/C/GHRS/1 6260 (Linear Extent and Ion Condition in = Lyman-Alpha Cloud: Cycle 5 Observations) The High Resolution Spectrograph (Side One) was used to make = spectrographic observations between 1215 A0 and 1500 A0 of the QSO = Q0107-025A in order to study the linear extent and the ion condition in = the Lyman-Alpha cloud. The binary QSO system will be searched for any = absorption systems which may not be in common with the two objects, thus = setting the limits on the linear size of its clouds. The observations = were completed as planned and no anomalies were reported. 1.6 Completed FOC/96 6931 (Exploratory Focus Monitoring Test) The Faint Object Camera (f/96) was used to make calibration images = as part of an exploratory focus monitoring test. An internal flat = calibration image was also taken. The observations were executed as = scheduled, and no anomalies were noted. 1.7 Completed Two Sets of FOS/BL 6913 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make a = series of stability calibration measurements of the wavelength location = and the location of the spectra. All the observations were internals. A = wavelength calibration image was also taken. The observations were = completed as planned, and no problems were reported. 1.8 Completed Two Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.9 Completed GHRS/2 6530 (The Wind Momentum: Luminosity Relationship = for LMC A and B Supergiants) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations the A/B class, supergiant star HD269905 which = is located in the Large Magellanic Cloud. These data are part of a large = effort to calibrate the wind momentum-luminosity function for supergiant = stars. The observations were completed as planned, and no problems were = reported. 1.10 Completed FOS/BL 6683 (Eclipsing Binaries in the Magellanic = Clouds - Fundamental Properties A) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of O/B Class, eclipsing binary system star = LMC-HV2274 located in the Large Magellanic Cloud. From these data and = the latest Kurucz model atmospheres, accurate temperatures will be = derived. The observations were executed as scheduled, and no anomalies = were noted. 1.11 Completed WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 1.12 Completed GHRS/2 6656 (Hydrogen, Interstellar Accretion, and Hot = Helium White Dwarfs) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the hot, helium-rich, DB class, White = Dwarf star PG0112+104. A theory predicts that DB stars with effective = temperatures between 30,000 deg and 45,000 deg K. masquerade as DA stars = due to a thin layer of hydrogen that floats on the surface. A = wavelength calibration image was also taken. The observations were = completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: There were no significant operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH